how do u work this problem out????????????

1. y=2x
6x-y=8

Found 2 solutions by Alan3354, checkley75:
Answer by Alan3354(21583)

y=2x
6x-y=8
-------
sub for y in the 2nd eqn
6x - 2x = 8
4x = 8
x = 2
------
y = 4

Answer by checkley75(3666)

y=2x
6x-y=8
substituting 2x for y in the second equation:
6x-(2x)=8
6x-2x=8
4x=8
x=8/4
x=2 ans.
y=2*2=4 ans.
Proof:
6*2-4=8
12-4=8
8=8